# Dynamical probe of the pseudo Jahn-Teller effect in one-dimensional confined fermions

### v1

@@ IV Dynamics in the case of weak confinement $\ell_{I}\approx\sqrt{N_{B}}\ell_{B}$
Figure 3: Spatiotemporal dependence of the one-body density of the bath $\rho_{B}^{(1)}(x;t)$ and impurity $\rho_{I}^{(1)}(x;t)$ species for an interacting system with $g=2.5\hbar\ell_{B}\omega_{B}$. The impurity mass $m_{I}=4m_{B}$ and trapping frequency $\omega_{I}=0.5\omega_{B}$ as well as the smaller shift $x_{s}=0.5\ell_{B}$ are considered. The different columns correspond to the distinct methods (see labels). For the impurity species, we include cyan dashed lines at $x=\pm(x_{s}+\sqrt{2}\ell_{I})$ as a visual reference to the oscillation amplitude in the non-interacting case.
@@ IV.1.1 Density analysis
Figure 4: Spatiotemporal dependence of the one-body density of the bath $\rho_{B}^{(1)}(x;t)$ and impurity $\rho_{I}^{(1)}(x;t)$ species for an interacting system with $g=5.0\hbar\ell_{B}\omega_{B}$. The impurity mass $m_{I}=4m_{B}$ and trapping frequency $\omega_{I}=0.5\omega_{B}$ as well as the smaller shift $x_{s}=0.5\ell_{B}$ are considered. The different columns correspond to the distinct methods (see labels). For the impurity species, we include cyan dashed lines at $x=\pm(x_{s}+\sqrt{2}\ell_{I})$ as a visual reference to the oscillation amplitude in the non-interacting case.
By introducing a weak interaction ($g=2.5\hbar\omega_{B}\ell_{B}$) between the bath and the single impurity, we present in Fig. 3 the corresponding time evolutions of the densities $\rho^{(1)}_{B}(x;t)$ and $\rho^{(1)}_{I}(x;t)$. Starting with the numerically exact results from the ML-X method, we observe a similar dipole oscillation of the impurity, see Fig. 3(b1), when compared to the non-interacting case, see Fig. 2(b1). However, the interaction introduces key differences from the non-interacting evolution. In particular, by comparing the interacting case of Fig. 3(a1) and (b1) to the non-interacting results of Fig. 2(a1) and (b1) we observe that due to the bath-impurity repulsion the bath species develops a density valley at the position of the impurity species. Additionally, in the time instances that the impurity crosses $x_{I}\approx 0$, e.g. for $t\approx 4\omega_{B}^{-1},12\omega_{B}^{-1}$, its density maximum decreases. In addition in the elapsed time between these transits an impurity density accumulation is observed for $x\approx-0.5$ for $104<\omega_{B}t<$ 10 and $x\approx 0.5$ for $1812<\omega_{B}t<$ 18, rendering the form of the impurity oscillation non sinusoidal in contrast to the non-interacting case, see Fig. 2. This effect is even more pronounced in the bath one-body density, leading to the pronounced formation of distinct minima in the spatial regions where the impurity accumulation occurs for $x_{B}\approx 0.5~{}\ell_{B}$ and $x_{B}\approx-0.5~{}\ell_{B}$, such that the bath depletion has a noticeably rectangular oscillation pattern. Another notable phenomenon for $\rho^{(1)}_{I}(x;t)$ is the reduction of its oscillation period, by comparing the oscillation pattern in Fig. 2(b1) with that in Fig. 3, we clearly see that the oscillation period is increased in the interacting case. Notice that both the ML-X and MCBO methods yield identical results, as both are numerically exact ab initio approaches (not shown here for brevity).
@@
To explore this possibility, we perform an one-body density analysis of the $g=5~{}\hbar\omega_{B}\ell_{B}$ case in Fig. 4. Starting again with the one-body density of the bath, Fig. 4(a1), and impurity species, Fig. 4(b1), within ML-X. Similar to Fig. 3(a1) and (b1), we observe density minima in the bath one-body density (a1) and again a minimum for $x_{I}\approx 0$ in the impurity one-body density. As expected, since these effects are a consequence of the interaction strength, it becomes more dominant for the strongly interacting case. In addition, the detected prolonging of the dipole oscillation period of the impurity species increases. An additional effect, hardly noticeable in the case of smaller interactions, compare Fig. 3(b1) to 4(b1), is that when the impurity species crosses the center of the harmonic trapping potential, for $g=5.0\hbar\omega_{B}\ell_{B}$ we detect a small part of the impurity species that is reflected at $x_{I}\approx 0$ and performs a mirror-symmetric oscillation. Furthermore, we observe that the impurity species dipole oscillation amplitude is significantly reduced when compared to the non-interacting case, compare Fig. 3(b1) to Fig. 2(b1). The oscillation curves deviate from the initial sinusoidal profile observed in the dipole oscillation for the non-interacting case Fig. 2(b1) and possess an approximately rectangular shape.
@@ IV.1.2 Analysis in terms of PEC
However, a deeper look at the population transfer oscillations reveals an open point in our analysis. In particular, inspecting Fig. 5(b) and 5(c), we observe that quite peculiarly the frequency of the population transfer does not match the frequency of the dipole oscillations of the impurity, see Fig. 3(b1) and 4(b1) respectively. More specifically, by inspecting Fig. 5(b), in the case of $g=2.5~{}\hbar\omega_{B}\ell_{B}$ we observe that the first minimum of $n_{1}(t)$ occurs at $t\approx 6.2\omega_{B}^{-1}$, where $\rho^{(1)}_{I}(x;t)$, see Fig. 3(b), is close to its minimum negative displacement, while the second maximum is at $t\approx 12.1\omega_{B}^{-1}$, when $\rho^{(1)}_{I}(x;t)$ crosses $x=0$. Finally, the second minimum visible in Fig. 3(b) corresponds to $t=16.6~{}\omega_{B}^{-1}$, where the impurity travels from its maximum positive displacement toward $x=0$. Similar patterns are also identified for the strongly interacting $g=5~{}\hbar\omega_{B}\ell_{B}$ state. It can then be claimed that there is little correlation of PEC transfer processes and the synthetic conical intersections and PJTE given the not straightforward relation of the PEC population transfer and the location where it occurs. However, by further analysis of the occupied impurity configurations within each PEC we can reveal the hidden relationship between these transfer processes and the PJTE.
@@
with its corresponding eigenfunctions, $\phi_{j,m}(x)$, and eigenenergies, $E_{j,m}$, shown in Fig. 7. For the non-interacting case and for both the ground-state PEC [Fig. 7(a1)], and the first excited PEC [Fig. 7(b1)], we identify an effective harmonic potential ($\sim x^{2}$), which is the same for both the lowest and the higher-lying PECs, which remain identical for both levels apart from a constant shift, reflecting the spatially uniform nature of the PECs in the non-interacting regime. Introducing interactions ($g=2.5$ in Fig. 7(a2) and 7(b2)) between the bath and impurity species results in a double-well potential for the lowest PEC, whereas the first excited state exhibits a triple-well structure. These features become more pronounced in the strongly interacting regime ($g=5.0\hbar\omega_{B}\ell_{B}$ Fig. 7(a3) and (b3)), with the triple-well potential of the $j=2$ PEC emerging particularly clearly. When examining our effective potential, we observe in Fig. 7(b2), (a3), and (b3) that, in addition to the triple-well structure, additional potential wells are present. However, these appear at such high energies that they do not significantly contribute to our analysis.
Figure 9: Spatiotemporal dependence of the one-body density of the bath $\rho_{B}^{(1)}(x;t)$ and impurity $\rho_{I}^{(1)}(x;t)$ species for an interacting system with $g=2.5\hbar\ell_{B}\omega_{B}$. The impurity mass $m_{I}=4m_{B}$ and trapping frequency $\omega_{I}=0.5\omega_{B}$ as well as the larger shift $x_{s}=1.5\ell_{B}$ are considered. The different columns correspond to the distinct methods (see labels). For the impurity species, we include cyan dashed lines at $x=\pm(x_{s}+\sqrt{2}\ell_{I})$ as a visual reference to the oscillation amplitude in the non-interacting case.
@@
details on the derivation of this expression is found in Appendix B. Considering the non-interacting case, Fig. 8(a) reveals that the impurity resides in a coherent state with populations consistent with $x_{s}=0.5\ell_{B}$. As previously mentioned the second PEC is completely depopulated in this case due to the absence of interactions, see Fig. 8(d). Introducing interactions with strength $g=2.5\hbar\omega_{B}\ell_{B}$ fundamentally alters the system’s behavior. By contrasting Fig. 8(b) to the non-interacting scenario shown in Fig. 8(a) and 8(d), we observe a shift of the initial population of the states of the first PEC, especially apparent is the reduction of the corresponding ground states for $m=0$. In addition, in contrast to the constant in time populations of Fig. 8(a), the occupations of the different $m$ contributions in the interacting case of Fig. 8(b), are not constant in time but they rather fluctuate with double the frequency of the impurity’s dipole oscillation. This behavior can be attributed to the impact of the Born-Huang term as it can be captured well by VABO (not shown here for brevity).
Figure 10: Spatiotemporal dependence of the one-body density of the bath $\rho_{B}^{(1)}(x;t)$ and impurity $\rho_{I}^{(1)}(x;t)$ species for an interacting system with $g=5.0\hbar\ell_{B}\omega_{B}$. The impurity mass $m_{I}=4m_{B}$ and trapping frequency $\omega_{I}=0.5\omega_{B}$ as well as the larger shift $x_{s}=1.5\ell_{B}$ are considered. The different columns correspond to the distinct methods (see labels). For the impurity species, we include cyan dashed lines at $x=\pm(x_{s}+\sqrt{2}\ell_{I})$ as a visual reference to the oscillation amplitude in the non-interacting case.
@@ IV.2.1 Density analysis
Analogously to Sec. IV.1.1, we begin by examining the weaker interaction, laying the groundwork for analyzing the bath and impurity one-body density before progressing to stronger coupling regimes. Hence, we consider in Fig. 9 the interaction strength $g=2.5\hbar\omega_{B}\ell_{B}$. Here, Fig. 9(a1) illustrates the one-body density of the bath, while Fig. 9(b1) presents the corresponding impurity one-body density, both derived from the numerically exact ML-X results. First, notice that as a result of the bath impurity interaction the initial state of the impurity for $g>0$ is displaced with respect to the $g=0$ case, compare the initial position of the density relative to the horizontal dashed lines in Fig. 9(b1) and Fig. 2(b2). The coherent dipole oscillation of the impurity species, evident in the non-interacting case for the complete time evolution, is only faintly visible up to $t\approx 7.5/\omega_{B}$ for $g=2.5~{}\hbar\omega_{B}\ell_{B}$. In the latter case it becomes apparent that a significant portion of the impurity density is unable to transmit through the bath and it is instead getting trapped within its spatial extent. This can be attributed to the scattering of the impurity with its bath leading to momentum transfer among the species. Beyond $t\approx 7.5\omega_{B}^{-1}$, $\rho^{(1)}_{I}(x;t)$ gets redistributed within its fermionic environment possessing substantial spatial modulations. In particular, it exhibits prominent interference-like peaks and troughs along the trajectory that a coherent state would have followed. While the bath density for $t<7.5\omega_{B}^{-1}$ shows a density depletion along the path of the coherent state, at later times it shows a highly fluctuating character, as shown in Fig. 9(a1). This behavior can also be attributed to the momentum exchange among the bath and impurity particles mentioned above.
@@
Next, we evaluate to what extent this behavior is captured by the adiabatic approximation. Focusing on the simpler NVABO approximation presented in Fig. 9(a2) and (b2), the dipole oscillation of the impurity species remains apparent in the impurity one-body density throughout the time evolution. The bath density exhibits a corresponding density depletion due to the bath-impurity repulsion, see Fig. 9(a2). In the impurity one-body density, a time-dependent modulation of the peak impurity density during its oscillation becomes noticeable after the first half-period, particularly visible in Fig. 9(b2) around $x_{I}\approx 0$ and $x_{I}\approx\pm 1\ell_{B}$. Moreover, after the first period of oscillation, a small fraction of the impurity density fails to penetrate back into the bath density, see Fig.  9(b2) for $t\approx 15\omega_{B}^{-1}$, leading to a small density fraction lying outside the fermionic environment.
@@
Turning our attention to the Born-Huang correction within the VABO approximation, Fig. 9(a3) shows a similar behavior for the bath one-body density to the NVABO approximation. However, in Fig. 9(b3), it is evident that the Born-Huang term prevents the impurity species from penetrating the bath species after the first half period, $t>7.5\omega_{B}^{-1}$. The pronounced dipole oscillation seen in the NVABO case gets distorted for larger times, as the impurity density fragments into smaller components that exhibit a fluctuating density pattern. Even in the VABO case, however, the fraction of the impurity density that deviates from the dipole oscillation pattern is much smaller when compared to the numerically exact result in Fig. 9(b1). In agreement to our previous arguments, the period of the dipole oscillation is consistent with the ML-X result of Fig. 2(b1) and significantly longer than NVABO, see Fig. 2(b2).
@@
Indeed, by turning our attention to the strongly interacting scenario ($g=5.0\hbar\omega_{B}\ell_{B}$) we observe a dramatic influence of non-adiabatic effects on the dynamics. Starting with the numerically exact result within ML-X, we observe significant deviations for both the bath, see Fig. 10(a1), and the impurity, see Fig. 10(b1), species compared to the previously analyzed cases. Unlike the lower interaction strength in Fig. 9 ($g=2.5\hbar\omega_{B}\ell_{B}$) or a smaller shift ($x_{s}=0.5\ell_{B}$) in Fig. 4, in the case of Fig. 10(b1) the impurity’s one-body density is largely unable to penetrate the bath species and it is entirely reflected, performing a small amplitude oscillation near the upper boundary of the bath density ($x_{I}\approx 2.5\ell_{B}$). This oscillation occurs locally and with a higher frequency, than the one set by $\omega_{I}$. During the oscillation dynamics, a small fraction of the impurity density can enter the bath density.
Consequently, the bath species exhibit a notably different behavior: it maintains its initial shape, characterized by a maximum density at the center, and does not develop pronounced density valleys in the one-body density due to the interaction. However, the bath species perform collective dipole oscillations in response to the impurity. Once a sizable fraction of the impurity density is able to penetrate into the bath density at $t\approx 10\omega_{B}^{-1}$, the latter starts exhibiting pronounced minima due to interaction.

### v2

@@ IV Dynamics in the case of weak confinement $\ell_{I}\approx\sqrt{N_{B}}\ell_{B}$
Figure 3: Spatiotemporal dependence of the one-body density of the bath $\rho_{B}^{(1)}(x;t)$ and impurity $\rho_{I}^{(1)}(x;t)$ species for an interacting system with $g=2.5\hbar\ell_{B}\omega_{B}$. The impurity mass $m_{I}=4m_{B}$ and trapping frequency $\omega_{I}=0.5\omega_{B}$ as well as the smaller shift $x_{s}=0.5\ell_{B}$ are considered. The different columns correspond to the distinct methods (see labels). For the impurity species, we include cyan dashed lines at $x=\pm(x_{s}+\sqrt{2}\ell_{I})$ as a visual reference to the oscillation amplitude in the non-interacting case. The dotted blue circles indicate the spatiotemporal region for which the impurity crosses $x_{I}=0$, corresponding to the location of an avoided crossing.
@@ IV.1.1 Density analysis
Figure 4: Spatiotemporal dependence of the one-body density of the bath $\rho_{B}^{(1)}(x;t)$ and impurity $\rho_{I}^{(1)}(x;t)$ species for an interacting system with $g=5.0\hbar\ell_{B}\omega_{B}$. The impurity mass $m_{I}=4m_{B}$ and trapping frequency $\omega_{I}=0.5\omega_{B}$ as well as the smaller shift $x_{s}=0.5\ell_{B}$ are considered. The different columns correspond to the distinct methods (see labels). For the impurity species, we include cyan dashed lines at $x=\pm(x_{s}+\sqrt{2}\ell_{I})$ as a visual reference to the oscillation amplitude in the non-interacting case. The dotted blue circles indicate the spatiotemporal regions where the impurity crosses $x_{I}=0$, corresponding to the location of an avoided crossing. The green dotted circles in (b1) highlight the reflected fraction of the impurity density by the avoided crossing at $x_{I}=0$.
By introducing a weak interaction ($g=2.5\hbar\omega_{B}\ell_{B}$) between the bath and the single impurity, we present in Fig. 3 the corresponding time evolutions of the densities $\rho^{(1)}_{B}(x;t)$ and $\rho^{(1)}_{I}(x;t)$. Starting with the numerically exact results from the ML-X method, we observe a similar dipole oscillation of the impurity, see Fig. 3(b1), when compared to the non-interacting case, see Fig. 2(b1). However, the interaction introduces key differences from the non-interacting evolution. In particular, by comparing the interacting case of Fig. 3(a1) and (b1) to the non-interacting results of Fig. 2(a1) and (b1) we observe that due to the bath-impurity repulsion the bath species develops a density valley at the position of the impurity species. Additionally, in the time instances that the impurity crosses $x_{I}\approx 0$, which are marked by the dotted blue circles, e.g. for $t\approx 4\omega_{B}^{-1},12\omega_{B}^{-1}$, its density maximum decreases. In addition in the elapsed time between these transits an impurity density accumulation is observed for $x\approx-0.5$ for $104<\omega_{B}t<$ 10 and $x\approx 0.5$ for $1812<\omega_{B}t<$ 18, rendering the form of the impurity oscillation non sinusoidal in contrast to the non-interacting case, see Fig. 2. This effect is even more pronounced in the bath one-body density, leading to the pronounced formation of distinct minima in the spatial regions where the impurity accumulation occurs for $x_{B}\approx 0.5~{}\ell_{B}$ and $x_{B}\approx-0.5~{}\ell_{B}$, such that the bath depletion has a noticeably rectangular oscillation pattern. Another notable phenomenon for $\rho^{(1)}_{I}(x;t)$ is the reduction of its oscillation period, by comparing the oscillation pattern in Fig. 2(b1) with that in Fig. 3, we clearly see that the oscillation period is increased in the interacting case. Notice that both the ML-X and MCBO methods yield identical results, as both are numerically exact ab initio approaches (not shown here for brevity).
@@
To explore this possibility, we perform an one-body density analysis of the $g=5~{}\hbar\omega_{B}\ell_{B}$ case in Fig. 4. Starting again with the one-body density of the bath, Fig. 4(a1), and impurity species, Fig. 4(b1), within ML-X. Similar to Fig. 3(a1) and (b1), we observe density minima in the bath one-body density (a1) and again a minimum for $x_{I}\approx 0$ (see the blue dotted circles) in the impurity one-body density. As expected, since these effects are a consequence of the interaction strength, it becomes more dominant for the strongly interacting case. In addition, the detected prolonging of the dipole oscillation period of the impurity species increases. An additional effect, hardly noticeable in the case of smaller interactions, compare Fig. 3(b1) to 4(b1), is that when the impurity species crosses the center of the harmonic trapping potential, for $g=5.0\hbar\omega_{B}\ell_{B}$ we detect a small part of the impurity species that is reflected at $x_{I}\approx 0$ and performs a mirror-symmetric oscillation as indicated by the green dotted circles. Furthermore, we observe that the impurity species dipole oscillation amplitude is significantly reduced when compared to the non-interacting case, compare Fig. 3(b1) to Fig. 2(b1). The oscillation curves deviate from the initial sinusoidal profile observed in the dipole oscillation for the non-interacting case Fig. 2(b1) and possess an approximately rectangular shape.
@@ IV.1.2 Analysis in terms of PEC
Figure 7: Impurity dynamics on the PECs are shown in the two regimes of impurity displacement. In both subfigures the Born-Huang term $V_{11}^{\rm ren}(x_{I})$, which scales with the interaction strength $g$, is shown with a constant vertical offset for better comparison with the PECs. The avoided crossings, marked by vertical lines at positions $x_{i}$ for $i=-2,\dots,2$, coincide with the maximum peaks of the Born-Huang term $V_{11}^{\rm ren}(x_{I})$, underscoring the importance of non-adiabatic coupling and interaction strength in shaping the impurity dynamics. (a) Small shift regime $x_{s}=0.5\ell_{B}$: The impurity exhibits quasi-adiabatic transport, except near the avoided crossing $x_{0}=0$, where density is exchanged between the lowest and first excited PEC. The PECs are derived for an interaction strength of $g=5.0\,\hbar\ell_{B}\omega_{B}$, impurity mass $m_{I}=4m_{B}$, and impurity trapping frequency $\omega_{I}=0.5\omega_{B}$. To highlight the avoided crossings, the PECs are scaled and shifted linearly. This adjustment enhances visual contrast without altering the underlying physical trends. The width of the faintly colored regions on the PECs indicates the relative contributions to each PEC. (b) Large shift regime $x_{s}=1.5\ell_{B}$: The impurity becomes self-trapped in an effective harmonic potential in the outer region of the lowest PEC, exhibiting quasi-diabatic behavior. As a result, only a small fraction of the impurity density can tunnel out of this potential.
@@
However, a deeper look at the population transfer oscillations reveals an open point in our analysis. In particular, inspecting Fig. 5(b) and 5(c), we observe that quite peculiarly the frequency of the population transfer does not match the frequency of the dipole oscillations of the impurity, see Fig. 3(b1) and 4(b1) respectively. More specifically, by inspecting Fig. 5(b), in the case of $g=2.5~{}\hbar\omega_{B}\ell_{B}$ we observe that the first minimum of $n_{1}(t)$ occurs at $t\approx 6.2\omega_{B}^{-1}$, where $\rho^{(1)}_{I}(x;t)$, see Fig. 3(b), is close to its minimum negative displacement, while the second maximum is at $t\approx 12.1\omega_{B}^{-1}$, when $\rho^{(1)}_{I}(x;t)$ crosses $x=0$. Finally, the second minimum visible in Fig. 3(b) corresponds to $t=16.6~{}\omega_{B}^{-1}$, where the impurity travels from its maximum positive displacement toward $x=0$. Similar patterns are also identified for the strongly interacting $g=5~{}\hbar\omega_{B}\ell_{B}$ state. It can then be claimed that there is little correlation of PEC transfer processes and the synthetic conical intersections and PJTE given the not straightforward relation of the PEC population transfer and the location where it occurs. To illustrate the discussed impurity dynamics, we provide a schematic representation in Fig. 7(a), focusing on the relevant PECs and the contribution of the Born-Huang term $V_{\rm ren}^{11}(x_{I})$. By applying a linear shift to the Born-Huang term, we can qualitatively compare its spatial behavior with the PECs and highlight that its maxima coincide with the avoided crossings. Additionally, the PECs are rescaled to emphasize the appearance of these avoided crossings. As previously derived, this illustrates how the impurity undergoes quasi-diabatic transport governed by the structure of the PECs, except near the avoided crossing at $x_{0}\approx 0$, where an exchange of contribution occurs between the ground and first excited states. However, by further analysis of the occupied impurity configurations within each PEC we can reveal the hidden relationship between these transfer processes and the PJTE.
@@
with its corresponding eigenfunctions, $\phi_{j,m}(x)$, and eigenenergies, $E_{j,m}$, shown in Fig. 8. For the non-interacting case and for both the ground-state PEC [Fig. 8(a1)], and the first excited PEC [Fig. 8(b1)], we identify an effective harmonic potential ($\sim x^{2}$), which is the same for both the lowest and the higher-lying PECs, which remain identical for both levels apart from a constant shift, reflecting the spatially uniform nature of the PECs in the non-interacting regime. Introducing interactions ($g=2.5$ in Fig. 8(a2) and 8(b2)) between the bath and impurity species results in a double-well potential for the lowest PEC, whereas the first excited state exhibits a triple-well structure. These features become more pronounced in the strongly interacting regime ($g=5.0\hbar\omega_{B}\ell_{B}$ Fig. 8(a3) and (b3)), with the triple-well potential of the $j=2$ PEC emerging particularly clearly. When examining our effective potential, we observe in Fig. 8(b2), (a3), and (b3) that, in addition to the triple-well structure, additional potential wells are present. However, these appear at such high energies that they do not significantly contribute to our analysis.
@@
details on the derivation of this expression is found in Appendix B. Considering the non-interacting case, Fig. 9(a) reveals that the impurity resides in a coherent state with populations consistent with $x_{s}=0.5\ell_{B}$. As previously mentioned the second PEC is completely depopulated in this case due to the absence of interactions, see Fig. 9(d). Introducing interactions with strength $g=2.5\hbar\omega_{B}\ell_{B}$ fundamentally alters the system’s behavior. By contrasting Fig. 9(b) to the non-interacting scenario shown in Fig. 9(a) and 9(d), we observe a shift of the initial population of the states of the first PEC, especially apparent is the reduction of the corresponding ground states for $m=0$. In addition, in contrast to the constant in time populations of Fig. 9(a), the occupations of the different $m$ contributions in the interacting case of Fig. 9(b), are not constant in time but they rather fluctuate with double the frequency of the impurity’s dipole oscillation. This behavior can be attributed to the impact of the Born-Huang term as it can be captured well by VABO (not shown here for brevity).
Figure 10: Spatiotemporal dependence of the one-body density of the bath $\rho_{B}^{(1)}(x;t)$ and impurity $\rho_{I}^{(1)}(x;t)$ species for an interacting system with $g=2.5\hbar\ell_{B}\omega_{B}$. The impurity mass $m_{I}=4m_{B}$ and trapping frequency $\omega_{I}=0.5\omega_{B}$ as well as the larger shift $x_{s}=1.5\ell_{B}$ are considered. The different columns correspond to the distinct methods (see labels). For the impurity species, we include cyan dashed lines at $x=\pm(x_{s}+\sqrt{2}\ell_{I})$ as a visual reference to the oscillation amplitude in the non-interacting case. The blue dotted line in the upper panels (ai) indicates the average position of the bath density $\langle x_{B}(t)\rangle$. The green dotted line in the impurity density panel (bi) gives as a guide to the eye the dipole oscillation in the NVABO case.
@@ IV.2 Role of interaction for the larger shift
Figure 11: Spatiotemporal dependence of the one-body density of the bath $\rho_{B}^{(1)}(x;t)$ and impurity $\rho_{I}^{(1)}(x;t)$ species for an interacting system with $g=5.0\hbar\ell_{B}\omega_{B}$. The impurity mass $m_{I}=4m_{B}$ and trapping frequency $\omega_{I}=0.5\omega_{B}$ as well as the larger shift $x_{s}=1.5\ell_{B}$ are considered. The different columns correspond to the distinct methods (see labels). For the impurity species, we include cyan dashed lines at $x=\pm(x_{s}+\sqrt{2}\ell_{I})$ as a visual reference to the oscillation amplitude in the non-interacting case. The blue dotted line in the upper panels (ai) indicates the average position of the bath density $\langle x_{B}(t)\rangle$. The green dotted line in the impurity density panel (bi) gives as a guide to the eye the dipole oscillation in the NVABO case.
@@ IV.2.1 Density analysis
Analogously to Sec. IV.1.1, we begin by examining the weaker interaction, laying the groundwork for analyzing the bath and impurity one-body density before progressing to stronger coupling regimes. Hence, we consider in Fig. 10 the interaction strength $g=2.5\hbar\omega_{B}\ell_{B}$. Here, Fig. 10(a1) illustrates the one-body density of the bath, while Fig. 10(b1) presents the corresponding impurity one-body density, both derived from the numerically exact ML-X results. First, notice that as a result of the bath impurity interaction the initial state of the impurity for $g>0$ is displaced with respect to the $g=0$ case, compare the initial position of the density relative to the horizontal dashed lines in Fig. 10(b1) and Fig. 2(b2). The coherent dipole oscillation of the impurity species, evident in the non-interacting case for the complete time evolution, is only faintly visible up to $t\approx 7.5/\omega_{B}$ for $g=2.5~{}\hbar\omega_{B}\ell_{B}$. In the latter case it becomes apparent that a significant portion of the impurity density is unable to transmit through the bath and it is instead getting trapped within its spatial extent. This can be attributed to the scattering of the impurity with its bath leading to momentum transfer among the species. Beyond $t\approx 7.5\omega_{B}^{-1}$, $\rho^{(1)}_{I}(x;t)$ gets redistributed within its fermionic environment possessing substantial spatial modulations. In particular, it exhibits prominent interference-like peaks and troughs along the trajectory that a coherent state would have followed. While the bath density for $t<7.5\omega_{B}^{-1}$ shows a density depletion along the path of the coherent state, at later times it shows a highly fluctuating character, as shown in Fig. 10(a1) and is represented by the oscillating average mean position of the bath species $\langle x_{B}(t)\rangle={\int\mathrm{d}x~{}x\rho_{B}^{(1)}(x,t)}/{\int\mathrm{d}x~{}\rho_{B}^{(1)}(x,t)}$, depicted by the dotted blue line. This behavior can also be attributed to the momentum exchange among the bath and impurity particles mentioned above.
Next, we evaluate to what extent this behavior is captured by the adiabatic approximation. Focusing on the simpler NVABO approximation presented in Fig. 10(a2) and (b2), the dipole oscillation of the impurity species remains apparent in the impurity one-body density throughout the time evolution. The bath density exhibits a corresponding density depletion due to the bath-impurity repulsion, see Fig. 10(a2). In the impurity one-body density, a time-dependent modulation of the peak impurity density during its oscillation becomes noticeable after the first half-period, particularly visible in Fig. 10(b2) around $x_{I}\approx 0$ and $x_{I}\approx\pm 1\ell_{B}$. Moreover, after the first period of oscillation, a small fraction of the impurity density fails to penetrate back into the bath density, see Fig.  10(b2) for $t\approx 15\omega_{B}^{-1}$, leading to a small density fraction lying outside the fermionic environment.
@@
Turning our attention to the Born-Huang correction within the VABO approximation, Fig. 10(a3) shows a similar behavior for the bath one-body density to the NVABO approximation. However, in Fig. 10(b3), it is evident that the Born-Huang term prevents the impurity species from penetrating the bath species after the first half period, $t>7.5\omega_{B}^{-1}$, as evidenced by the loss of agreement with the dipole oscillation predicted by the NVABO approximation. The pronounced dipole oscillation seen in the NVABO case gets distorted for larger times, as the impurity density fragments into smaller components that exhibit a fluctuating density pattern. Even in the VABO case, however, the fraction of the impurity density that deviates from the dipole oscillation pattern is much smaller when compared to the numerically exact result in Fig. 10(b1). In agreement to our previous arguments, the period of the dipole oscillation is consistent with the ML-X result of Fig. 2(b1) and significantly longer than NVABO, see Fig. 2(b2).
@@
Indeed, by turning our attention to the strongly interacting scenario ($g=5.0\hbar\omega_{B}\ell_{B}$) we observe a dramatic influence of non-adiabatic effects on the dynamics. Starting with the numerically exact result within ML-X, we observe significant deviations for both the bath, see Fig. 11(a1), and the impurity, see Fig. 11(b1), species compared to the previously analyzed cases. Unlike the lower interaction strength in Fig. 10 ($g=2.5\hbar\omega_{B}\ell_{B}$) or a smaller shift ($x_{s}=0.5\ell_{B}$) in Fig. 4, in the case of Fig. 11(b1) the impurity’s one-body density is largely unable to penetrate the bath species and it is entirely reflected, performing a small amplitude oscillation near the upper boundary of the bath density ($x_{I}\approx 2.5\ell_{B}$). This oscillation occurs locally and with a higher frequency, than the one set by $\omega_{I}$ as evidenced by the green dotted line. During the oscillation dynamics, a small fraction of the impurity density can enter the bath density.
Consequently, the bath species exhibit a notably different behavior: it maintains its initial shape, characterized by a maximum density at the center, and does not develop pronounced density valleys in the one-body density due to the interaction. However, the bath species perform collective dipole oscillations in response to the impurity, as indicated by the time evolution of the average mean position shown in the plot. Once a sizable fraction of the impurity density is able to penetrate into the bath density at $t\approx 10\omega_{B}^{-1}$, the latter starts exhibiting pronounced minima due to interaction.